OpenStudy (anonymous):

what is the solution to the rational equation x/2x+1- 1/4=2/2x+1? Please help

OpenStudy (anonymous):

\[\frac{x}{2x-1}-\frac{1}{4}=\frac{2}{2x+1}\] You can multiply both sides by 2x+1 \[x-\frac{1}{4}(2x+1)=2\] From then on it should be routine. Remember that the solution cannot be x=-1/2 because that makes the denominator 0.

OpenStudy (anonymous):

Sorry typo on my first equation, though I'm sure you'll understand anyway. :D

OpenStudy (anonymous):

Yeah, thanks so much!